OpenStudy (anonymous):

If the length of each side of a cube is doubled, is the surface area doubled? Use an example to support your answer. @cp9454

OpenStudy (cp9454):

no, new surface area becomes 4 times of the previous one.

OpenStudy (kirbykirby):

An easy example: Take the side to be 1 cm. Then the surface area is 6*(1*1)=6 If you double the side, the side has length 2 cm. Then the surface area is 6*(2*2) = ?

OpenStudy (cp9454):

let a cube of edge a cm

OpenStudy (cp9454):

then its surface area will be \[6a^2\]

OpenStudy (cp9454):

now the side is doubled i.e. 2a, so its surface area will be \[6(2a)^2 = 6*4a^2\]